OpenStudy (anonymous):

BAYSIDE INSURANCE OFFERS TWO HEALTHPLANS A, GISELLE WOULD HAVE TO PAY THE FIRST $130 PLUS 40% OF THE REST. UNDER PLAN B GISELLE WOULD PAY THE $200 BUT ONLY 25% OF THE REST. FOR WHAT AMOUNT OF MEDICAL BILLS WILL PLAN B SAVE GISELLA MONEY. ASSUME SHE HAS OVER $200 IN BILLS

OpenStudy (anonymous):

A=0.4X + 130 B= 0.25X+200

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

If I deserve medal then give me one.

OpenStudy (anonymous):

WELL I NEED THE ANSWER FIRST PLEASE BUT YES I GIVE MEDAL FOR HELPING ME!!!!

OpenStudy (anonymous):

what question?

OpenStudy (anonymous):

you plug x for amout you want

OpenStudy (anonymous):

GISELLE WOULD SAVE WITH PLAN B IF SHE HAD MORE THAN $ ?? IN BILLS?

OpenStudy (anonymous):

Fot plan B if the nill is $ 200 B= 0.25(200)+200 B = 50+200 B= $250

OpenStudy (anonymous):

sorry the bill not nill I tye wrong

OpenStudy (anonymous):

is the bill more more than $ 200 you use the formula to plug in and solve

OpenStudy (anonymous):

any question? I be back

OpenStudy (anonymous):

SO IS THAT THE ANSWER THEN NANCY...

OpenStudy (anonymous):

not yet I'm working

OpenStudy (anonymous):

0.4X + 130= 0.25X+200 0.4X - 0.25X=200 -130 0.15X=70 X = 130/0.15 X = $866.7 The bill $866.7 or more

OpenStudy (anonymous):

TWO ANGLES ARE COMPLEMENTARY. THE SUM OF THE MEASURE OF THE FIRST ANGLE AND ONE-FOURTH THE SECOND ANGLE 77.55 DEGREE. FIND THE MEASURE OF THE ANGLES? WHAT IS THE MEASURE OF THE SMALLER ANGLE?

OpenStudy (anonymous):

what is picture?

OpenStudy (anonymous):

ALSO WHAT IS MEASURE OF THE ANGLE?

OpenStudy (anonymous):

NOT A PICTURE I JUST NEED THE ANSWERS PLEASE

OpenStudy (anonymous):

can you check made sure not missing, not engough information

OpenStudy (anonymous):

2 angle but doesn't saided vertical angle aor adcent angle

OpenStudy (anonymous):

FIRST ANGLE AND ONE-FOURTH deosn't said regtanor square, trap,roh I don't know ,if you have more detail I may know

OpenStudy (anonymous):

I think 51.22

OpenStudy (anonymous):

180-77.55=102.45 102.45/2=51.23 you can post this question , some one may know better than me